COLOR CENTER LASER SPECTROSCOPY OF THE ν_{1} BAND OF DCN-HCN GENERATED IN A PULSED SUPERSONIC JET EXPANSION

\underline{K. HARADA}, M. ISHIGURO, M. SHIRASAKA, K. TANAKA and T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashi-ku, Fukuoka 812-81, Japan.

Color center laser spectroscopy has been applied to direct absorption observation of the ν_{1} band of DCN-HCN generated in a pulsed supersonic jet expansion. A new wavenumber scanning system has been used for the observation, where a laser end mirror and an internal etalon have been locked to a fringe of an external scanning etalon and the wavenumber was swept by scanning the external etalon. The 15 lines were assigned to the ν_{1} band of DCN-HCN. In order to obtain precise molecular constants, we have measured direct absorption millimeter wave spectra of deuterated species of HCN dimer in the ground state. A simultaneous analysis of the infrared and millimeter wave spectra yielded the band origin and rotational constants of DCN-HCN as ν_{1}=3241.60809(93) cm^{−1}, B_{1}=1666.66(47) MHz, and B_{0}=1662.4350(35) MHz.